Question 1: what is the yield to maturity of a corporate bond with 20 years to maturity, a coupon rate of 6% per year, a $1000 par value and current market price of $1100? assume semiannual coupon payments?

Answer 4.27%; 4.76%; 5.01%; 5.19%; 5.35%
Answer #1

Coupon rate(semi-annual) = 6%/2 = 3%

Coupon = 1000*3% = 30

Number of periods = 20*2 = 40

semi annual yield = r

1100 = 30/(1+r)+30/(1+r)^2+.....+30/(1+r)^39+30/(1+r)^40+1000/(1+r)^40 [market price of bond is sum of all the cash flows]

by trial and error method we can see value of r = 2.595%

So, YTM = 2.595*2 = 5.19% (Ans)

With r = 2.595%

Period Cashflow Discounted cashflow
1 30 29.24119109
2 30 28.50157521
3 30 27.78066691
4 30 27.07799299
5 30 26.39309225
6 30 25.72551513
7 30 25.07482346
8 30 24.44059015
9 30 23.82239889
10 30 23.21984394
11 30 22.6325298
12 30 22.06007095
13 30 21.50209168
14 30 20.95822572
15 30 20.42811611
16 30 19.91141489
17 30 19.40778292
18 30 18.91688964
19 30 18.43841282
20 30 17.97203843
21 30 17.51746033
22 30 17.07438017
23 30 16.64250711
24 30 16.22155768
25 30 15.8112556
26 30 15.41133155
27 30 15.02152302
28 30 14.64157418
29 30 14.27123561
30 30 13.91026425
31 30 13.55842317
32 30 13.21548143
33 30 12.88121393
34 30 12.55540126
35 30 12.23782959
36 30 11.92829045
37 30 11.62658068
38 30 11.33250225
39 30 11.04586213
40 1030 369.6488779
Total 1100.058815